Grade 12th passGeneral Physics

A body has started rest and is accelerated uniformly. The distance covered by body in first 3s is equal to the distance travelled by it in the last second. Find the time of journey of body

Jaya Lakshmi Jujjuri
10 Years agoGrade 12th pass

1 Answer

surudhi
10 Years ago
The distance covered in 3s:
s = ut + ½ at2 = ½ at(since the body was at rest u = 0)
s = ½ a(9) = 9/2 a
The distance travelled in last second:
s= a/2 (2n – 1) where n denotes the particular second
Since,
9 (a/2) = a/2 ( 2n – 1)
9 = 2n – 1
10 = 2n
5 = n
 
The total time taken is 5s.